The width of a rectangle is 3 feet less than its length. If the perimeter is 22 feet, what is the width? (Pre-Algebra) By Charles P. Mckeague, fifth addition.

Answer by nerdybill(7384)

Let x = length
then from "width of a rectangle is 3 feet less than its length" we have
x-3 = width
.
Perimeter = 2(length+width)
22 = 2(x + x-3)
22 = 2(2x-3)
22 = 4x-6
28 = 4x
7 feet = x (length)
.
width:
x-3 = 7-3 = 4 feet (width)